FOURIER TRANSFORM MICROWAVE SPECTRA OF N_{2}-(CH_{3})_{2}O

\underline{YOSHIYUKI KAWASHIMA}, YOSHIO TATAMITANI, YASUMASA MORITA, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN; and EIZI HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

As an extension of the studies on the dynamical behavior of van der Waals complexes such as those on CO-DME, we have investigated nitrogen - dimethyl ether complex N_{2}-(CH_{3})_{2}O, by using Fourier transform microwave spectroscopy. We have scanned the frequency region from 6 to 25 GHz and have found four sets of a-type rotational transitions ranging from J = 2 ← 1 up to J = 6 ← 5 for N_{2}-DME and ^{15}N_{2}-DME and two sets for ^{15}NN-DME. Two of the four sets (referred to as group I) of N_{2}-DME and ^{15}N_{2}-DME have large centrifugal distortion constants. Each rotational transition of N_{2}- DME showed complicated splitting patterns due to the quadrupole coupling of the two nitrogen atoms and the number of hyperfine components was much smaller for group I than for the other (group II). This observation indicates that the group I complexes involve para-N_{2} and the group II ortho-N_{2}. In the case of ^{15}NN-DME only one type (corresponding to group II) of the complexes was detected because of the lack of symmetry. Some of the a-type transitions observed for 15N_{2}-DME consisted of closely spaced triplets; the splittings, which were nearly independent of the quantum numbers J, were ascribed to the internal rotation of the two methyl tops of DME. The observed transition frequencies of N_{2}-DME, 15N_{2}-DME, and ^{15}NN-DME were analyzed for each set separately, by using an ordinary asymmetric-rotor Hamiltonian. The inertial defects I_{cc} − I_{aa} − I_{bb} thus obtained for N_{2}-DME were -29.31 (10) and -30.97 (10) uÅ^{2} for the two sets of group I and -9.98 (9) and -12.58 (11) uÅ^{2} for group II. These results indicated that the heavy-atom skeleton of N_{2}-DME was not planar. The observed moments of inertia were analyzed to give the distance between the centers of gravity of the two component molecules, DME and N_{2}, to be approximately 3.45 Å. By assuming a Lennard-Jones-type potential the dissociation energy was estimated to be E_{B} = 0.74 ∼ 1.17 kJ mol^{−1}, to be compared with the values 1.0 and 2.5 kJ mol^{−1} for Ne-DME and Ar-DME, respectively. MP2/6-31++g(d, p) calculations suggest that N_{2}-DME is non-planar and is bound by a very flat potential energy surface, in qualitative agreement with our results.